ANNES AGE IS SEVEN-TENTHS OF BILLS AGE.IN 4 YEARS ANNES AGE WILL BE EIGHT-ELEVENTHS OF BILLS AGE.HOW OLD IS BILL NOW?
Answer by richwmiller(17219)

Stop shouting!
All caps is shouting!
a=7/10*b
a+4=8/11(b+4)
a=28
bill =40 years
check
28=7/10*40
28=28
32=8/11(44)
32=32
ok